OpenStudy (amtran_bus):

Find an equation of the tangent to the curve at the point by first eliminating the parameter

OpenStudy (amtran_bus):

x=e^t , y=(t-6)^2 ; (1,36)

OpenStudy (amtran_bus):

I tried to elem. the parameter and got lnx^2 - 12lnx +36 ?

OpenStudy (amistre64):

why would we want to eliminate the parameter? just because it says so?

OpenStudy (amtran_bus):

Correct.

OpenStudy (amtran_bus):

Did I even do it right?

OpenStudy (amistre64):

ln(x) = t is suitable sure

OpenStudy (amtran_bus):

Hum. So what do I do to get the equation? I cant just plug in (1,36) for x and y because ln(1) is 0.

OpenStudy (amistre64):

well, since y = (t-6)^2, and t=ln(x) y = (ln(x)-6)^2

OpenStudy (amistre64):

in order to make a line, we need a point and a slope ... how do we find the slope?

OpenStudy (amtran_bus):

or lnx^2-12lnx+36

OpenStudy (amistre64):

no need to expand it ..

OpenStudy (amtran_bus):

Hum. Ok. So y-y1=m(x-x1)

OpenStudy (amistre64):

good, and we have y1 and x2, we only need m

OpenStudy (amistre64):

typos lol

OpenStudy (amtran_bus):

So just take delta y/delta x? Which is ... um...

OpenStudy (amistre64):

which is chain rule ...

OpenStudy (amtran_bus):

Oh. Duh. Gotta take deriv. of that.

OpenStudy (amistre64):

yep

OpenStudy (amtran_bus):

Give me a sec!

OpenStudy (amtran_bus):

2(lnx-6)*1/x

OpenStudy (amistre64):

very good, and when x=1? m=?

OpenStudy (amtran_bus):

-6?

OpenStudy (amistre64):

2(0-6)/1 is not -6

OpenStudy (amtran_bus):

It is if you are lazy and plug it in your calc wrong. :)

OpenStudy (amistre64):

lol

OpenStudy (amtran_bus):

-12

OpenStudy (amistre64):

so whats our line?

OpenStudy (amtran_bus):

y=-12x+

OpenStudy (amtran_bus):

gott get that last number.

OpenStudy (amistre64):

y-36 = -12(x-1) is the equation of a line ... but format however you want

OpenStudy (amtran_bus):

Okie dokie. So just -12x+1+36? So y=-12x+37?

OpenStudy (amistre64):

-12x -12(-1) +36 so no

OpenStudy (amtran_bus):

Story of my life. How about -12x+12+36 or y=-12x+48

OpenStudy (amistre64):

looking good lol

OpenStudy (amtran_bus):

So algebra is harder than calculus. Thank you.

OpenStudy (amistre64):

i agree :)